# Robust Multimodal Learning with Missing Modalities via Parameter-Efficient Adaptation

### v3

@@ 4.5 Experiments for Multimodal Sentiment Analysis
We tested our adaptation method for multimodal sentiment analysis on CMU-MOSI [69] and CMU-MOSEI [3] datasets, and report the results in Table 5. We use multimodal transformer (MulT) [53] as the base model and adapt it using our approach. We observed that when text is available and either audio or video or both are missing at the test time, the performance does not drop significantly. Similar trend was reported in [18]. If text is missing at test time, then the performance of the Pretrained model drops significantly. The Adapted models can partially compensate for missing text and offer significantly better performance compared to the pretrained model.
@@
For CMU-MOSI dataset, we observe 1.69% and 2.44% improvement in accuracy over the Pretrained model performance when only audio and only visual are available, respectively. In all these scenarios, we also see larger improvement in F1 score. The adapted model offers significant improvement when audio-visual modalities are available and text is missing. It shows 6.56% improvement in terms of accuracy and 12.01% improvement in terms of F1 score over the Pretrained model. For CMU-MOSEI dataset, we see even greater improvement in all the metrics. Experiments show 25.7%, 24.21% and 21.41% improvement in accuracy for audio only, visual only and audio-visual scenarios compared to the Pretrained model. We also observe 27.91%-36.30% improvement in F1 score compared to the pretrained model.
@@
We also compare our adaptation method with existing methods for multimodal sentiment analysis. The table shows that for CMU-MOSI dataset, BERT_MAG works better in terms of accuracy but our adaptation method works better in terms of F1 score. One thing to mention is that BERT_MAG uses a pretrained BERT model and finetunes it on the dataset but we are not using any pretraining on extra data. For CMU-MOSEI, our adaptation method works better for most of the cases. Which also demonstrates the generalizability and applicability of our approach to diverse modalities and different tasks.
@@ 5 Conclusion
Missing modalities at test time can cause significant degradation in the performance of multimodal systems. In this paper, we presented a simple and parameter-efficient adaptation method for robust multimodal learning with missing modalities. We demonstrated that simple linear operations can efficiently transform a single pretrained multimodal network and achieve performance comparable to multiple (independent) dedicated networks trained for different modality combinations. We evaluated the performance of our method and compared with existing robust methods for different multimodal segmentation and sentiment analysis tasks. Our method requires an extremely small number of additional parameters (e.g., $<0.7\%$ of the total parameters in most experiments), while significantly improving performance compared to existing baseline models and methods for different missing modality scenarios. Our adaptation strategy is applicable to different network architectures and tasks, which can be a versatile solution to build robust multimodal systems. In this work, our main focus was to enhance robustness of existing multimodal models and we did not try to optimize the number of parameters or find the optimal places to insert those learnable layers. Future study will explore these areas to further reduce the number of parameters and application of the approach beyond the tasks and datasets used in this study.
Reproducibility Statement
We are committed to ensuring the reproducibility of our research and to facilitate the broader scientific community in replicating and building upon our work. The source code and trained models are available at github [^1] https://github.com/CSIPlab/Robust-multimodal-learning. We have provided a clear and comprehensive README.md file to guide users in setting up the environment, running the code, and reproducing the results in the paper. We outline the specific data preprocessing steps, list of hyperparameters and configurations used in our experiments in Section 4.2 in the main text and Section 2 in the supplementary section. We hope this makes it easy for others to replicate our experiments. We have provided scripts and instructions in our source code to reproduce the main experimental results presented in this paper. Additionally, we have provided pretrained models allowing others to directly reproduce the results.
Ethics Statement
To the best of our knowledge this work does not give rise to any significant ethical concerns.
@@ 1 Datasets
Table 6: Hyperparameters for the experiments on CMU-MOSI and CMU-MOSEI datasets for multimodal sentiment analysis.
@@
2 Implementation Details
MFNet Dataset: We divied the 4 channel RGB-T images into three channel RGB and one channel thermal images. Then data pre-processing and augmentation was applied following CMNeXt from [70]. MiT-B4 from [62] was the backbone for the base CMNeXt model. One set of scale and shift parameters was learnt for each input modality combination. Input images were sized at $640\times 480$ for both training and testing and we report single scale performance for all the experiments. The scale and shift parameters were trained for 100 epochs with a batch size of 4.
CMU-MOSI and CMU-MOSEI Datasets: We used Multimodal Transformer (MulT) from [53] as the base model. Preprocessed datasets and all the configurations are available on the repository [^7] https://github.com/thuiar/MMSA. First we trained the multimodal transformer (MulT) model on all the available modalities and then adapted the pretrained model for different missing modality scenarios. The hyperparameters for the experiments are shown in Table 6.
Table 7: Learnable parameter counts for different parameter efficient model adaptation methods. As seen from the table, scale and shift introduce less than 0.7% of the total model parameters.
Method Total Parameters (M) Learnable Parameters (M) % of Total Parameters
@@ 2 Implementation Details
Table 8: Performance comparison with parameter efficient model adaptation techniques on CMNeXt model for MFNet dataset. Average column indicates average performance when one of the two modalities gets missing. Mean accuracy, F1 score and % mIoU are shown for all the experiments.
@@
Table 9: Performance comparison with parameter efficient model adaptation techniques on CMNeXt model for NYUDv2 dataset. Average column indicates average performance when one of the two modalities gets missing. Mean accuracy, F1 score and % mIoU are shown for all the experiments.
@@
Table 10: Performance comparison with different parameter efficient model adaptation techniques on CMNeXt model for MCubeS dataset. Average column indicates the average performance. Mean accuracy, F1 score and % mIoU are shown for all the experiments.
@@
Table 11: Per class % IoU comparison between pretrained and adapted CMNeXt model on MFNet and MCubeS datasets. Adapted model show better performance for most of the classes leading to overall performance improvement. Here A, D and N stand for Angle of Linear Polarization (AoLP), Degree of Linear Polarization (DoLP) and Near-Infrared (NIR) respectively.
(a) Per class % IoU of pretrained and adapted CMNeXt model on MFNet dataset.
@@
(b) Per class % IoU of pretrained and adapted CMNeXt model on MCubeS dataset.
@@
3 Number of Learnable Parameters
We report the number of learnable parameters for different parameter-efficient adaptation techniques (for multimodal segmentation) in Table 7. We insert scale and shift layers after each linear, convolutional and norm (both batch norm and layer norm) layers. The number of learnable parameter varies with the size of the backbone. We used MiT-B4 as the backbone while counting these learnable parameters. Scale and shift adds only 0.789M learnable parameters which is less than 0.7% of the total model parameters. Despite this very few parameters, it improves performance significantly in different missing modality scenarios. For this study we mainly focused on improving missing modality robustness and did not try to optimize the number of learnable parameters. We will leave that part for future studies.
Figure 3: Visualization of predicted segmentation maps for pretrained and adapted models on MFNet and NYUDv2 datasets for multimodal semantic segmentation and MCubeS dataset for multimodal material segmentation. Only RGB images are shown from MCubeS dataset for brevity. CMNeXt column shows the predictions when all the modalities are available. Segmentation quality improves significantly after model adaptation for all the input modality combinations. A, D and N stand for angle of linear polarization, degree of linear polarization and near-infrared respectively.
@@ 3 Number of Learnable Parameters
4 Performance Comparison with Parameter Efficient Model Adaption Techniques
We performed a detailed performance comparison with other parameter efficient methods for the three segmentation datasets. The results are summarized in Table 8 for RGB-thermal segmentation on MFNet dataset, Table 9 for RGB-depth segmentation on NYUDv2 dataset and Table 10 for multimodal material segmentation on MCubeS dataset. For each method, we take a model trained on all the available modalities. Then we freeze the pretrained weights and tune the learnable parameters for the corresponding adaption method. We have shown mean accuracy, F1 score and % mIoU for each experiment.
4.1 Performance Comparison for RGB-Thermal Semantic Segmentation on MFNet Dataset
Table 8 summarizes the results on MFNet dataset when the base CMNeXt model is adapted with other parameter efficient model adaptation techniques. Experiments show that scale and shift shows the best performance in all three matrices compared to all other methods. It shows a significant improvement of +8.46% in mIoU, +9.33% in F1 score and +17.48% in mean accuracy on an average over the pretrained model. The average performance is also close to dedicatedly trained models.
4.2 Performance Comparison for RGB-Depth Semantic Segmentation on NYUDv2 Dataset
Similar trend is observed for RGB-Depth semantic segmentation on NYUDv2 dataset as shown in Table 9. Scale only and BitFit adapted models show slightly better performance for some of the matrices. But in most of the cases scale and shift adapted model performs better. For all the matrices, scale and shift shows a significant improvement of +16.54% in mIoU, +22.05% in F1 score and +23.99% in mean accuracy over the pretrained model on an average and consistently outperforms dedicated training.
4.3 Performance Comparison for Multimodal Material Segmentation on MCubeS Dataset
@@
We show comparison with parameter efficient model adaptation techniques on MCubeS dataset in Table 10. Scale and shift outperforms all other methods in most of the matrices for all input combinations. It also shows an improvement of +3.99% in mIoU, +3.57% in F1 score and +3.74% in mean accuracy on an average over the pretrained model. Furthermore, Scale and shift also outperforms dedicated training for all input modality combinations. These experiments corroborate the fact that scale and shift provides better model adaption for different missing modality scenarios.
5 Per Class IoU Comparison
@@
To further analyze how the adaption is helping the model improve overall semantic and material segmentation performance, we conduct a per-class % intersection over union (IoU) analysis on the pretrained and adapted models. Table 11(a) and 11(b) summarize the results. We show the per class % IoU comparison for different missing modality situations on MFNet dataset on Table 11(a). From the table we can see that when RGB is available and thermal is missing, the adaptation helps improve performance for most of the classes. Though we see some performance drop for bike (-0.74%) and guardrail (-5.25%) classes, the rest of the classes have better % IoU than the pretrained model. Bump (+6.26%), person (+4.42%), and curve (+4.14%) classes show greater improvement after adaptation. When thermal is available and RGB is missing, adaptation improves performance for all the classes. Among the classes, bike (+39.76%), car (+27.59%), car stop (+24.04%), color cone (+19.37%) and bump (+13.91%) are showing impressive performance improvement over the pretrained model.
Results for MCubeS dataset is shown on Table 11(b). Here A, D, and N stand for angle of linear polarization (AoLP), degree of linear polarization (DoLP) and near-infrared (NIR) respectively. Experiments show that when only RGB is available and the rest of the modalities are missing, the adapted model performs better in detecting all the 20 classes present in the dataset. Gravel (36.2%), asphalt (16.1%), rubber (15.2%), wood (12.9%) and sky (10.2%) are some of the classes who show the most performance boost after adaptation. In other input combinations, most of the classes see performance improvement compared to the pretrained model. Though we see some performance drop in a few classes, most of the classes show improvement in % IoU which leads to the overall performance improvement after adaption.
6 Visualization of Predicted Segmentation Maps

### v4

@@ 4.4 Experiments on Multimodal Sentiment Analysis
We tested our adaptation method for multimodal sentiment analysis on CMU-MOSI [58] and CMU-MOSEI [59] datasets, and report the results in Table 4. We use multimodal transformer (MulT) [62] as the base model and adapt it using our approach. We observed that when text is available and either audio or video or both are missing at the test time, the performance does not drop significantly. Similar trend was reported in [15]. If text is missing at test time, then the performance of the base MulT model drops significantly. The Adapted models can partially compensate for missing modality and offer significantly better performance compared to the base MulT model.
@@
For CMU-MOSI dataset, we observe 1.69% and 2.44% improvement in accuracy and larger improvement in F1 score over the base MulT model when only audio and only visual are available, respectively. The adapted model offers significant improvement when audio-visual modalities are available and text is missing. It shows 6.56% improvement in accuracy and 12.01% improvement in F1 score over the base MulT model. For CMU-MOSEI dataset, we see even greater improvement in all the metrics. Experiments show 25.7%, 24.21% and 21.41% improvement in accuracy for audio only, visual only and audio-visual scenarios compared to the MulT model. We also observe 27.91%-36.30% improvement in F1 score compared to the base MulT model.
@@
We compare our adaptation method with existing methods for multimodal sentiment analysis. For CMU-MOSI dataset, BERT_MAG works better in terms of accuracy but our adaptation method works better in terms of F1 score. One thing to mention is that BERT_MAG uses a pretrained BERT model and finetunes it on the dataset but we are not using any pretraining on extra data. For CMU-MOSEI, our adaptation method works better for most of the cases.
[…]
Available Modality ViLT Attention Input Adapted
Image Text [64] Prompts [27] Prompts [27] (Ours)
100% 30% 66.29 72.57 74.53 75.38
30% 100% 76.66 86.05 86.18 88.31
65% 65% 69.25 78.09 79.08 81.77
100% 0% 63.60 67.70 68.10 67.66
0% 100% 76.10 85.30 84.80 86.01
Average Accuracy 70.38 77.94 78.54 79.83
Total Params (M) 112.26 112.49 112.49 112.47
Learnable Params (M) 0.0 0.221 0.221 0.207
Change (%) +0.0% +0.20% +0.20% +0.18%
4.5 Experiments on Multimodal Action Recognition
We evaluate our approach on NTU RGB+D [57] dataset for multimodal action recognition task. We use UMDR [63] as the base model and adapt it using our approach. As shown in Table 5, our adaptation performs better compared to recent modality masking and generation based approach ActionMAE [23] and modality de- and re-coupling based approaches Motion-RGBD [91] and UMDR [63]. Our adaptation shows 7.03% and 1.06% improvement over ActionMAE and UMDR respectively when RGB is available and depth is missing. We see 3.79% and 0.30% improvement over ActionMAE and UMDR respectively when depth is available and RGB is missing. Moreover, our method outperforms all the existing baseline methods in all the scenarios. Which also indicates that our approach can learn better feature representation compared to modality masking, generation and distillation based approaches.
The base UMDR model has 75.82M parameters. Our adaptation method adds 0.24M additional learnable parameters, which is only 0.32% of the total model parameters. Other methods in this table do not report their total parameter counts, so we omit the total parameters column for this table.
4.6 Experiments on Multimodal Classification
To further evaluate the effectiveness of our approach, we compare it with recent prompt based approach missing-aware prompts [27] on UPMC Food-101 [60] dataset. The results are summarized in Table 6. For fair evaluation, we use the same experimental setup and evaluation script as [27] to evaluate performance on different available and missing modality scenarios. Image and text columns indicate the amount of image and text modality available during both training and testing. Our adaptation method outperforms prompting based approach for most of the scenarios. On an average, our approach shows 1.29% improvement over the best prompting method and 9.45% improvement over the base ViLT model. These results corroborate the fact that adapting models by intermediate feature modulation helps the model learn optimal feature representation to perform better on different missing modality scenarios.
Efficiency on Parameters. We keep the pre-trained ViLT backbone frozen and compare the additional learnable parameters required for the learnable prompts [27] and our method. We require less additional parameters while performing better than both input level and attention level prompts. Thus our adaptation method shows greater parameter efficiency and effectiveness compared to prompt based approach.
Figure 3: Cosine similarity between complete and missing modality features of the pretrained model (Pretrained) and complete and missing modality features of the adapted model (Adapted) on MCubeS and NTU RGB+D datasets. Adapted models show higher similarity to the complete modality features compared to the pretrained model, indicating less deviation and better handling of missing modalities.
4.7 Feature and Parameter Analysis
We perform additional analysis to evaluate the effectiveness and generalizability of the adaptation approach. We discuss them in the section.
4.7.1 Why adapted model performs better?
To further analyze how the adaption is helping the model improve overall performance, we conducted cosine similarity analysis of the final fused features extracted from the last layer of the network. To be specific, we calculate the cosine similarity between the complete and missing modality features from the pretrained model (Pretrained), and the cosine similarity between the complete and missing modality features from the adapted model (Adapted). We show the cosine similarities for each class in Figure 3.
The adapted model demonstrates a higher cosine similarity to the complete modality features compared to the pretrained model on both MCubeS dataset for multimodal material segmentation and NTU RGB+D dataset for multimodal action recognition when RGB is available and other modalities are missing. This increased similarity indicates that the adapted model better retains the essential information from the original complete modality input features, even when some modalities are missing. Consequently, this robustness in feature representation leads to a significant improvement in the model’s overall performance. These results demonstrate the effectiveness of the adapted model in handling scenarios with missing modalities and maintaining robust prediction quality.
We only show first 20 out of 60 classes for NTU RGB+D dataset here. We have included comparison for all the 60 classes including other missing scenarios in Section S6 in the supplementary materials.
4.7.2 Performance gain vs learnable parameters
Our method achieves significant performance gains with a small number of additional learnable parameters. As shown in Table 2 and 3, adapted models provide 8.46% and 16.54% improvement in mIoU on an average over the base CMNeXt model with only 0.79M additional parameters (i.e., 0.68% of the total model parameters). For multimodal sentiment analysis, as shown in Table 4, adapted models provide 3.57% and 23.78% improvement in accuracy and 6.79% and 33.22% improvement in F1 score for CMU-MOSI and CMU-MOSEI datasets, respectively over the base MulT model with only 0.02M additional parameters (i.e., 0.775% of the total model parameters). For multimodal classification on UPMC Food-101 dataset, as shown in Table 6, adapted models achieve an average performance improvement of 9.45% over the base ViLT model with only 0.207M additional learnable parameters (i.e., 0.18% of the total model parameters).
In summary, learning a small number of additional parameters in a base network provides significant performance improvement in the case of missing modalities across all tasks and architectures in our experiments. The parameter complexity of our approach is comparable/better than existing robust methods like CRM [20] in Table 2 and prompts [27] in Table 6. However, existing works on missing modality robustness vary widely in terms of model architectures [32, 4], fusion methods [35, 63], training procedures [20, 32], and missing feature generation methods [23]. Due to this heterogeneity, a fair comparison based solely on model size/number of parameters is infeasible.
5 Limitations and Future Directions
In this work, our main focus was to enhance missing modality robustness of existing multimodal models. Though our method can make existing models robust to different missing modality scenarios, it has certain limitations. First, we only considered missing modality during test time. However in real life scenarios, modalities can be missing in both train and test times. Second, we need to learn one set of adaptation parameters for every combination of missing modalities. While the number of adaptation parameters is small, the overall complexity can increase as the number of modality combinations increase. If we expect one modality to be missing at test time, adaptation parameter complexity would scale with the number of modalities. Third, we insert the learnable layers after each layer of the encoders and the fusion block. We did not try to optimize the number of parameters or find the optimal places to insert those learnable layers. Future study will explore these areas to further reduce the number of parameters, enhance the effectiveness and applicability of the approach in newer tasks and datasets.
@@ 6 Conclusion
Missing modalities at test time can cause significant degradation in the performance of multimodal systems. In this paper, we presented a simple and parameter-efficient adaptation method for robust multimodal learning with missing modalities. We demonstrated that simple linear operations can efficiently transform a single pretrained multimodal network and achieve performance comparable to multiple (independent) dedicated networks trained for different modality combinations. We evaluated the performance of our method and compared with existing robust methods for five different multimodal tasks. Our method requires an extremely small number of additional parameters (e.g., $<1\%$ of the total parameters in most experiments), while significantly improving performance compared to existing baseline models and methods for different missing modality scenarios. Our adaptation strategy is applicable to different network architectures, modalities and tasks, which can be a versatile solution to build robust multimodal systems.
@@ Acknowledgements
NTU RGB+D (NTU60) dataset [57] is a popular multimodal action recognition dataset. The dataset contains 56,880 action samples divied into 60 classes. The actions can be braodly categorized into three different categories: daily actions, medical conditions and mutual actions. It has four different input modalities: RGB videos $(1920\times 1080)$, depth map sequences $(512\times 424)$, infrared (IR) videos $(512\times 424)$ and 3D skeletal data (25 major body joints). Three Microsoft Kinect V2 cameras were used to capture the videos simultaneously. It has two evaluation protocols: cross subject and cross view. We used RGB and depth data for our experiments and evaluated on cross subject protocol.
@@
UPMC Food-101 dataset [60] is a popular challenging multimodal classification dataset. It has 90,704 image-text pairs divided into train, validation and test sets. The dataset is annotated for 101 classes. Classes are identical to the ETHZ Food-101 dataset [92]. The samples are noisy as they were collected in an uncontrolled environment and thus huge diversity among samples is observed.
S2 Implementation Details
MFNet Dataset: We divided the 4 channel RGB-T images into three channel RGB and one channel thermal images. Then data pre-processing and augmentation was applied following CMNeXt from [61]. MiT-B4 from [75] was the backbone for the base CMNeXt model. One set of scale and shift parameters was learnt for each input modality combination. Input images were sized at $640\times 480$ for both training and testing and we report single scale performance for all the experiments. The scale and shift parameters were trained for 100 epochs with a batch size of 4.
Table S1: Hyperparameters for the experiments on CMU-MOSI and CMU-MOSEI datasets for multimodal sentiment analysis.
@@
Table S2: Learnable parameter counts for different parameter efficient model adaptation methods. As seen from the table, scale and shift introduce less than 0.7% of the total model parameters.
Method Total Learnable % of Total
Parameters (M) Parameters Parameters
@@
Table S3: Performance comparison (mIoU) of different parameter-efficient adaptation techniques for MFNet, NYUDv2, and MCubeS datasets. Each column reports mIoU of the Adapted model with the corresponding modalities, and Avg indicates average performance. A and D denote Angle and Degree of Linear Polarization.
Datasets MFNet NYUDv2 MCubeS
Methods RGB Thermal Avg RGB Depth Avg RGB RGB-A RGB-A-D Avg
Pretrained 53.71 35.48 44.60 51.19 5.26 28.23 42.32 48.81 49.06 46.73
Dedicated 55.86 53.34 54.60 52.18 33.49 42.84 48.16 48.42 49.48 48.69
Scale Only 54.77 49.23 52.00 53.04 36.12 44.58 50.16 50.55 51.13 50.61
Shift Only 54.57 48.96 51.77 53.04 36.25 44.65 50.13 50.40 50.86 50.46
BitFit 54.39 49.07 51.73 53.09 36.64 44.87 50.19 50.57 51.07 50.61
LoRA 54.19 47.45 50.82 52.87 34.97 43.92 49.59 50.07 50.80 50.15
Norm 54.65 47.49 51.07 53.05 34.73 43.49 49.95 50.51 51.07 50.51
Table S4: Performance comparison with parameter efficient model adaptation techniques on CMNeXt model for MFNet dataset. Average column indicates average performance when one of the two modalities gets missing. Mean accuracy, F1 score and % mIoU are shown for all the experiments.
@@
NTU RGB+D (NTU60) Dataset: We followed the same pre-processing and experimental setup as [91, 63]. We used RGB and depth data for our experiments and evaluated our method using cross subject protocol for fair comparison with [88, 23, 91, 63]. We extract 16 frames per video following previous methods and utilize DSN and DTN for spatial and temporal information encoding following [91, 63]. Then we train the SSF layers for 20 epochs with a batch size of 6.
CMU-MOSI and CMU-MOSEI Datasets: We used Multimodal Transformer (MulT) from [62] as the base model. Preprocessed datasets and all the configurations are available on the repository [^6] https://github.com/thuiar/MMSA. First we trained the multimodal transformer (MulT) model on all the available modalities and then adapted the pretrained model for different missing modality scenarios. The hyperparameters for the experiments are shown in Table S1.
UPMC Food-101 Dataset: For multimodal classification on this dataset, we use ViLT as the base model and follow the experimental setup used by [27]. We use the same hyper-parameters and script for generating and evaluating different missing modality combinations. The SSF layers were trained for 10 epochs with a learning rate of $1e^{-5}$.
S3 Number of Learnable Parameters
We report the number of learnable parameters for different parameter-efficient adaptation techniques (for multimodal segmentation) in Table S2. We insert scale and shift layers after each linear, convolutional and norm (both batch norm and layer norm) layers. The number of learnable parameter varies with the size of the backbone. We used MiT-B4 as the backbone while counting these learnable parameters. Scale and shift adds only 0.789M learnable parameters which is less than 0.7% of the total model parameters. Despite this very few parameters, it improves performance significantly in different missing modality scenarios. For this study we mainly focused on improving missing modality robustness and did not try to optimize the number of learnable parameters. We will leave that part for future studies.
S4 Performance Comparison with Parameter Efficient Model Adaption Techniques
We performed a detailed performance comparison with other parameter efficient model adaptation methods for the three segmentation datasets. Comparison among different parameter-efficient adaptation methods show that SSF-based adaptation provides overall best performance. We summarize the results for scale only, shift only, BitFit [51], norm layer fine-tuning and LoRA [43] in Table S3. We also show detailed comparison in Table S4 for RGB-thermal segmentation on MFNet dataset, Table S5 for RGB-depth segmentation on NYUDv2 dataset and Table S6 for multimodal material segmentation on MCubeS dataset. For each method, we take a model trained on all the available modalities. Then we freeze the pretrained weights and tune the learnable parameters for the corresponding adaption method. We have shown mean accuracy, F1 score and % mIoU for each experiment.
Table S5: Performance comparison with parameter efficient model adaptation techniques on CMNeXt model for NYUDv2 dataset. Average column indicates average performance when one of the two modalities gets missing. Mean accuracy, F1 score and % mIoU are shown for all the experiments.
@@
Table S6: Performance comparison with different parameter efficient model adaptation techniques on CMNeXt model for MCubeS dataset. Average column indicates the average performance. Mean accuracy, F1 score and % mIoU are shown for all the experiments.
@@
Table S7: Per class % IoU comparison between pretrained and adapted CMNeXt model on MFNet dataset. Adapted model show better performance for most of the classes leading to overall performance improvement.
@@
Table S8: Per class % IoU comparison between pretrained and adapted CMNeXt model on MCubeS dataset. Adapted model show better performance for most of the classes leading to overall performance improvement. Here A, D and N stand for Angle of Linear Polarization (AoLP), Degree of Linear Polarization (DoLP) and Near-Infrared (NIR) respectively.
@@
Figure S1: Cosine similarity between complete and missing modality features of the pretrained model (Pretrained) and complete and missing modality features of the adapted model (Adapted) under different missing modality scenarios on MCubeS dataset. The adapted model shows higher similarity to the complete modality features compared to the pretrained model, indicating less deviation and better handling of missing modalities.
Figure S2: Cosine similarity between complete and missing modality features of the pretrained model (Pretrained) and complete and missing modality features of the adapted model (Adapted) under different missing modality scenarios on NTU RGB+D dataset. Adapted models show comparable/higher similarity to the complete modality features compared to the pretrained model, indicating less deviation and better handling of missing modalities.
@@
S4.1 Performance comparison with other parameter-efficient model adaption techniques
@@
Apart from robust models, we also compare different parameter-efficient adaptation techniques. We summarize the results in Table S3. For MFNet dataset, SSF outperforms all the methods and performance is significantly better than the Pretrained model and close to the Dedicated models. For NYUDv2 and MCubeS datasets, the Adapted model performs better than both Pretrained and Dedicated models. These experiments also show that SSF performs better than other methods for most of the input modality combinations for all the datasets. We show a detailed comparison for each dataset in terms of mean accuracy, F1 score and % mIoU in Table S4 - S6.
S4.2 Performance Comparison for RGB-Thermal Semantic Segmentation on MFNet Dataset
Table S4 summarizes the results on MFNet dataset when the base CMNeXt model is adapted with other parameter efficient model adaptation techniques. Experiments show that scale and shift shows the best performance in all three matrices compared to all other methods. It shows a significant improvement of +8.46% in mIoU, +9.33% in F1 score and +17.48% in mean accuracy on an average over the pretrained model. The average performance is also close to dedicatedly trained models.
S4.3 Performance Comparison for RGB-Depth Semantic Segmentation on NYUDv2 Dataset
Similar trend is observed for RGB-Depth semantic segmentation on NYUDv2 dataset as shown in Table S5. Scale only and BitFit adapted models show slightly better performance for some of the matrices. But in most of the cases scale and shift adapted model performs better. For all the matrices, scale and shift shows a significant improvement of +16.54% in mIoU, +22.05% in F1 score and +23.99% in mean accuracy over the pretrained model on an average and consistently outperforms dedicated training.
S4.4 Performance Comparison for Multimodal Material Segmentation on MCubeS Dataset
We show comparison with parameter efficient model adaptation techniques on MCubeS dataset in Table S6. Scale and shift outperforms all other methods in most of the matrices for all input combinations. It also shows an improvement of +3.99% in mIoU, +3.57% in F1 score and +3.74% in mean accuracy on an average over the pretrained model. Furthermore, Scale and shift also outperforms dedicated training for all input modality combinations. These experiments corroborate the fact that scale and shift provides better model adaption for different missing modality scenarios.
S5 Per Class IoU Comparison
@@
To further analyze how the adaption is helping the model improve overall semantic and material segmentation performance, we conduct a per-class % intersection over union (IoU) analysis on the pretrained and adapted models. Table S7 and S8 summarize the results. We show the per class % IoU comparison for different missing modality situations on MFNet dataset on Table S7. From the table we can see that when RGB is available and thermal is missing, the adaptation helps improve performance for most of the classes. Though we see some performance drop for bike (-0.74%) and guardrail (-5.25%) classes, the rest of the classes have better % IoU than the pretrained model. Bump (+6.26%), person (+4.42%), and curve (+4.14%) classes show greater improvement after adaptation. When thermal is available and RGB is missing, adaptation improves performance for all the classes. Among the classes, bike (+39.76%), car (+27.59%), car stop (+24.04%), color cone (+19.37%) and bump (+13.91%) are showing impressive performance improvement over the pretrained model.
Table S9: Per-class IoU comparison among Dedicated, Pretrained, and Adapted model on MFNet Dataset.
RGB Thermal
Class Dedicated Pretrained Adapted Dedicated Pretrained Adapted
Unlabeled 97.95 97.79 97.79 97.85 95.97 97.45
Car 88.21 87.62 88.06 85.84 55.24 83.39
Person 62.97 51.13 55.55 71.00 68.47 69.96
Bike 63.03 61.94 61.20 56.50 9.27 49.19
Curve 36.02 30.05 34.19 40.03 31.85 40.53
Car_Stop 40.45 39.36 40.52 25.70 2.75 25.14
Guardrail 11.35 21.04 15.78 7.52 0.00 1.55
Color_Cone 52.99 45.55 48.67 42.24 16.87 33.85
Bump 49.78 48.95 55.21 53.35 38.92 53.22
Average 55.86 53.71 55.22 53.34 35.48 50.48
Table S10: Per-class IoU comparison among Dedicated, Pretrained, and Adapted model on MCubeS Dataset.
RGB RGB+AoLP RGB+AoLP+DoLP
Class Dedicated Pretrained Adapted Dedicated Pretrained Adapted Dedicated Pretrained Adapted
Asphalt 85.68 69.74 85.80 87.45 83.20 84.43 87.02 84.45 84.14
Concrete 43.41 39.18 43.72 45.29 43.28 45.36 43.82 41.20 45.57
Metal 51.36 47.57 52.64 53.11 50.72 53.75 50.65 46.68 54.08
Road_Marking 64.95 67.33 73.79 59.53 72.63 74.50 71.29 72.76 74.58
Fabric 30.08 26.88 27.86 30.52 26.37 30.42 29.86 25.20 30.50
Glass 51.20 44.56 50.95 53.25 51.91 53.22 50.88 51.59 54.23
Plaster 0.11 0.16 0.75 0.37 0.21 0.63 0.37 0.33 0.63
Plastic 21.81 20.86 24.18 23.22 28.05 26.90 22.05 26.05 28.10
Rubber 26.11 15.24 30.40 27.96 22.23 28.76 27.63 28.76 30.15
Sand 59.65 61.76 67.80 58.75 67.67 68.97 62.88 66.72 69.01
Gravel 64.24 36.73 72.86 66.24 63.40 69.25 71.08 65.63 67.63
Ceramic 27.11 19.12 27.10 30.11 22.70 24.83 30.30 25.97 25.86
Cobblestone 68.76 67.15 68.06 71.14 67.48 67.51 71.25 66.47 67.80
Brick 41.34 35.96 42.93 41.90 40.55 43.24 43.52 40.38 43.79
Grass 58.93 49.46 57.57 56.89 54.44 58.36 56.77 49.92 58.01
Wood 45.12 36.13 49.02 44.99 44.91 48.21 44.64 45.08 49.07
Leaf 76.62 71.55 74.91 75.70 73.90 75.05 75.60 72.67 75.03
Water 45.12 36.13 43.41 41.46 44.77 48.13 52.39 49.43 53.73
Human 4.27 14.70 18.27 3.32 21.79 14.43 1.33 25.59 13.71
Sky 96.39 86.25 96.47 96.53 96.04 96.42 96.34 96.33 96.49
Average 48.11 42.32 50.42 48.39 48.81 50.62 49.48 49.06 51.11